Operations Report – April 4th

SOL: 4
Name of person filing report: Louis Joseph
Non-nominal systems: Suit 2 (ventilation valve), Suit 5 (battery) and Suit 9 (visor)
Notes on non-nominal systems: Louis and Hippolyte are going to have a look at Suit 5 battery on Sol 5 (tomorrow). One of the ventilation valves of Suit 2 is stuck. Arnaud and Louis tried to work on it but were unsuccessful. They are going to look more into it on Sol 5 as they are both not planned for an EVA.
Spirit rover used:
Hours: (before EVA) 250.1
Beginning charge: 100%
Ending charge: 82%
Currently charging: Yes (10:00 am)
Opportunity rover used:
Hours: (before EVA) 176.6
Beginning charge: 100%
Ending charge: 87%
Currently charging: On return from EVA, we didn’t plugged it as requested because of cloudy skies
Curiosity rover used:
Hours: (before EVA) 275
Beginning charge: 100%
Ending charge: 87%
Currently charging: On return from EVA, we didn’t plugged it as requested because of cloudy skies
Perseverance rover used:
Hours: 289.5
Beginning charge: 93%
Ending charge: 84%
Currently charging: We plugged it at 9:51
Hours: 289.5
Beginning charge: 84%
Ending charge: 78%
Currently charging: na
General notes on rovers: Perseverance’s handbrake seems to be a little loose. The rover is still going backwards even if the handbrake is on.
Summary of Hab operations: put notes here
WATER USE: 24.51 gallons
Water (static tank): 432.37 gallons
Static tank pipe heater (on or off):
Static tank heater (On or off)
Toilet tank emptied: no or yes
Summary of internet: Used for music playing, mails, research for experiments, learning languages, watching science videos.
Summary of suits and radios:
For the morning EVA:
Radios 1,3,4,6 were used.
Light Suits 1,3,4,6 were used.
For the afternoon EVA:
Radios 1,2,3,4,6 were used.
Light Suits 1,2,3,4,6 were used.
Summary of GreenHab operations: I watered the crops with 4.5 gallons of water and harvested some tomatoes for the evening meal and mint to make tea.
WATER USE: 4.5 Gallons
Heater: On
Supplemental light: N/A
Tomatoes (18 pcs) : 63 g
Mint (~40 leaves) : 3 g
Summary of ScienceDome operations:
During the afternoon (2pm to 3pm), Imane, Alba and Arnaud used the science dome for blood and saliva collection to assess stress and immunity markers as well as processing the blood samples after their collection. We have frozen the saliva samples for further analysis to come when the crew will be back to Belgium.
Between 3pm to 4pm, Hippolyte asked each of us to complete a questionnaire about our experience in evaluating an artificial voice intelligence system.
Dual split: Off
Summary of RAM operations: First of all, we used the RAM to work on a stuck valve for the suit number 2 using WD40. We also used the RAM for the second part of Romain’s experiment as in the first day. He conducted part of his TapStrap experiment.
Summary of any observatory issues: NA
Summary of health and safety issues: One person cut his finger a little while working with a pair of scissors. This was disinfected and a plaster applied. 1 person had a headache and one paracetamol 1g was given. He is now better.
Questions, concerns and requests to Mission Support: how is our water consumption looking?

Copyright © The Mars Society. All rights reserved. | Main Site